OpenStudy (anonymous):

1. A man divides P10,000 in two investments, one at 10% and the other at 30%. Find how much is invested at each rate so that the two investments produce the same income annually.

terenzreignz (terenzreignz):

seems you're back...

OpenStudy (anonymous):

HAHA .. o'yeah. :))

terenzreignz (terenzreignz):

Are you supposed to use systems here?

OpenStudy (anonymous):

yes..

terenzreignz (terenzreignz):

x and y... suppose he divided the 10000 into two portions, x and y...

terenzreignz (terenzreignz):

So we know for sure that x + y = 10000 yes?

OpenStudy (anonymous):

ehe? we will use x and y??

terenzreignz (terenzreignz):

Well, what do you want? a and b? p and q? You said use systems...

OpenStudy (anonymous):

ahh.. haha! :)) nvm,, i'm wrong..

terenzreignz (terenzreignz):

huh? So... do we continue?

OpenStudy (anonymous):

yes.. continue..

terenzreignz (terenzreignz):

x + y = 10000 we need another equation... how to find one?

OpenStudy (anonymous):

ahm,, by using the 2 investments??

terenzreignz (terenzreignz):

mhmm... so, the two investments produce the same income annually...

OpenStudy (anonymous):

aha! then?

terenzreignz (terenzreignz):

What's the income for the x?

OpenStudy (anonymous):

the amount invested at 10%?

terenzreignz (terenzreignz):

Whichever, it doesn't really matter... suppose x was the amount invested at 10 % then...

OpenStudy (anonymous):

hm.. i don't know the next..

terenzreignz (terenzreignz):

How much does x earn after a year? 10% of x correct? So 0.1x

terenzreignz (terenzreignz):

the amount y is invested at 30% so how much does it earn?

OpenStudy (anonymous):

10000-0.1x?

terenzreignz (terenzreignz):

no... :/ 0.3y

OpenStudy (anonymous):

ah.. i got it.. :)) sorry.. :(

terenzreignz (terenzreignz):

So, they have the same income, so you have 0.1x = 0.3y along with x + y = 10000 now solve the system :P

OpenStudy (anonymous):

just substitute???

terenzreignz (terenzreignz):

Whatever works... for 0.1x = 0.3y you can ease it up by multiplying both sides by 10 x = 3y now it's simpler. Now solve...

OpenStudy (anonymous):

y= 2500 and x=7500 .. ? right?

terenzreignz (terenzreignz):

there you go.

OpenStudy (anonymous):

eh, how about if we will use 1 variable only on that problem??

terenzreignz (terenzreignz):

Well, just downright divide the 10000 into two parts, x and the rest, which is 10000-x. And you know that 0.1(x) = 0.3(10000-x) you'll arrive at the same answer...

OpenStudy (anonymous):

ahh.. owkie. :) thank you. :)